how many uniform identical textbooks with a width of 25cm can be stacked on top of each other on a level of surface without the stack falling over,if each successive book is displaced 3 cm in width direction relative to next lower book..?

Grade:

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Farsi

25/3 = 8+ so 8 + the bottom book = 9

if the pile is 9 high, the center of gravity is still inside the bottom book.
and more so for all the books above that.

the 10th book will make the pile fall.

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..